The fill amount of bottles of a soft drink is normally distributed, with a mean of 2.0 liters and a standard deviation of 0.05 liter. If you select a random sample of 25 bottles, what is the probability that the sample mean will be
(a) between 1.99 and 2.0 liters?

Answer by stanbon(75887)

The fill amount of bottles of a soft drink is normally distributed, with a mean of 2.0 liters and a standard deviation of 0.05 liter. If you select a random sample of 25 bottles, what is the probability that the sample mean will be
(a) between 1.99 and 2.0 liters?
z(1.99) = (1.99-2)/[0.05/sqrt(25)] = -0.01/0.01 = -1
z(2.0) = 0
-------
P(1.99< x-bar < 2) = normalcdf(-1,0) = 0.34
-----------------------------
Cheers,
Stan H.
=====================